# How Do AI Agents Do Human Work? Comparing AI and Human Workflows Across Diverse Occupations

### v1

@@ 7 Related Work
Agent Performance at Work  With the rise of computer-use agents, most efforts have focused on non-work settings such as online shopping (Yao et al., 2022; Zhou et al., 2024) and travel planning (Deng et al., 2023). Beyond acting as personal assistants, agents hold growing potential to automate professional work and are projected to exert substantial labor market impacts (Grace et al., 2018; Eloundou et al., 2023). Several studies have examined agents’ autonomous performance in specific domains, including engineering (Jimenez et al., 2024), software services (Drouin et al., 2024), design (Si et al., 2025b; Ge et al., 2025), research (Wijk et al., 2024; Lu et al., 2024; Si et al., 2025a), and general business activities (Wornow et al., 2024; Shen et al., 2024; Huang et al., 2025); as well as their collaboration with humans in legal (Choi and Schwarcz, 2024; Das et al., 2025) and writing (Clark et al., 2018) contexts. TheAgentCompany benchmark (Xu et al., 2024) advances this line of inquiry by evaluating multiple occupations, though it remains confined to the scope of a software engineering company. In this work, we seek to achieve a holistic coverage of all computer-related occupations, estimated to affect 71.9% of daily job functions, providing a more representative view of agent performance at work.
@@
Understanding Human Workflows  Prior research has examined human workflows in certain professional activities, such as customer service (Brynjolfsson et al., 2025b), design (Son et al., 2024a), and writing (Dang et al., 2025). More recently, with the rise of AI agents, many studies have explored how human workflows change when AI tools are introduced into work. These studies observe a shift in human activities from ‘doing’ to ‘managing’, which tends to reduce cognitive load (Shukla et al., 2025), yet may also decrease overall productivity (Shukla et al., 2025; Simkute et al., 2025) and promote dishonest behaviors (Köbis et al., 2025).
@@
While these studies offer insights on how humans use and adapt to AI agents, most overlook the scenario in which AI agents automate human work entirely — arguably the most concerning case for the workforce. In such automation settings, human workflows differ substantially from those in augmentation scenarios (Chen et al., 2025). As employment for entry-level positions already faces threats (Brynjolfsson et al., 2025a), a head-to-head comparative analysis between agents and humans — covering both behavior patterns and outcome quality — is crucial for identifying human advantages and informing policy and design decisions. Existing comparative efforts have limited contexts, such as ambiguity resolution in web tasks (Son et al., 2024b), a single web design task (loo, 2025), or secondary analysis on research articles (Vaccaro et al., 2024). To truly understand how agents as workers may reshape today’s human-centered workspace, we argue for a fine-grained, workflow-level investigation spanning multiple occupations and skill domains, to enable findings that shed light on diverse areas of work.
@@
Inducing Computer Workflows  Several tools have been developed for recording human trajectories (Wang et al., 2025b; Shaikh et al., 2025). However, such raw activity data are often hard to interpret, as low-level actions are task-agnostic and lack clear semantic grounding. Even for humans themselves, articulating explicit goals for certain actions, e.g., random clicks on blank areas driven by hesitation or anxiety, can be challenging. This poses challenges in conducting the head-to-head comparisons between human and agent computer-use activities. Inspired by how humans naturally communicate and coordinate around higher-level goals, we conceptualize and model activities at a coarser granularity through workflows, defined as sequences of actions frequently undertaken to achieve a specific goal (Wang et al., 2025c). Prior efforts toward this goal have largely depended on manual annotation (Sodhi et al., 2023; Grunde-McLaughlin et al., 2025). In contrast, we introduce the first automated tool capable of inducing interpretable workflows from raw mouse and keyboard traces, generating structured goals at multiple levels of granularity to support diverse analytical needs. While we primarily employ this tool to study agent and human worker behaviors in this work, it also holds promise for broader applications, such as agent skill learning and human-agent collaboration.
@@ 8 Conclusion
The O*NET Database  The O*NET database provides a holistic view of occupational tasks across the U.S. labor market. However, the inherent difficulty in constructing and maintaining such a large-scale database may introduce inaccuracies. Although we utilized data from the most recent 2024 release, it may not fully reflect the current workforce distribution. Task requirements for certain occupations may have evolved, while others may have become obsolete or newly emerged in the current job market (Acemoglu et al., 2022).
Understanding AI’s Impact on Work  It is natural for many to wonder “will AI take my job?” especially amid the rapid advances in AI over the past few years. Beyond offering insights to researchers of AI agents and human-AI interactions, we hope this work also speaks to the broader community of human workers whose occupations may be affected by AI adoption. Because many in the general public may have little exposure to how AI actually functions, we aim for this study to serve as an accessible and detailed account that helps demystify their capabilities and limitations — encouraging informed awareness rather than ungrounded fear or hype. While parts of our study may reveal the uncomfortable reality that agents may gradually take over fixed-procedure tasks as they advance, many essential aspects of human work, such as communication and decision-making, still demand substantial technological progress and regulatory considerations before comparable automation becomes feasible (Comunale, 2024). Overall, our goal is to offer a clearer, evidence-based understanding of how these seemingly powerful AI agents operate and to inform a grounded, proactive perspective on how humans and AI can move forward together.
@@ Acknowledgments
We would like to thank Sherry Tongshuang Wu, Yuxuan Li, Christina Qianou Ma, Zhoujun Cheng, Xuhui Zhou, Yuxiao Qu, Lawrence Jang, Xinran Zhao, Saujas Vaduguru, Vijay Viswanathan, Chenyang Yang, Mingqian Zheng, Youssouf Kebe, John Yang, Vishakh Padmakumar, Jiaxin Pei, and Alexander Spangher for their helpful discussions and insightful feedback on the project. Zora Zhiruo Wang is supported by Google PhD Fellowship.
@@ A.3 Human Activity Collection
Processing Raw Computer-Use Activities  Computer-use activities collected from human real-world usage can be redundant, due to the limitations in the recording tool implementation. For example, each mouse scrolling (scroll()) and keyboard press (keypress(‘h’)) event is recorded as an independent action. These human actions execute with arguments that are more fine-grained (e.g., keypress(‘h’), keypress(‘o’), $\cdots$, keypress(‘?’)) than arguments in the agent actions (e.g., keypress(‘how ai agents do human work?’)), thus becoming costly to process and hard to interpret. We hence perform post-processing to merge consecutive keypress and scroll actions into single actions with concatenated arguments. We also identify possible double clicks if two click() actions happen within 0.1 seconds; and merge them into a single double_click() action. This effectively simplifies the trajectory data, reducing the action count by 83.2% (from 5831.1 to 981.1 actions) in an average human trajectory, meanwhile aligning better with the granularity of agent computer activities (Wang et al., 2025b).
@@ A.4 Agent Activity Collection
While we can directly access and store all actions and states for OpenHands agents, we can only access the information presented at the ChatGPT and Manus agent UI interfaces. For each step in the ChatGPT and Manus trajectories, the agent will elaborate its thought process in text, and present a screenshot of the virtual computer screen. We infer the actions taken from the agent’s thoughts by mapping back to the common computer-use agent action space such as Zhou et al. (2024); Xie et al. (2024). We present the actions employed by all agent frameworks in Table 4. Nonetheless, action names for ChatGPT and Manus may not be precisely what was executed by the agent, but are our best guesses and should serve a similar functionality to the actual action, which we unfortunately do not have access to. Meanwhile, the listed actions are only actions that we observe agents have taken while completing all the task instances in our study. These actions may not constitute the full action space available to the agents.
@@ C.1 Do Agents Follow Human Task Workflows?
As shown in Table 7, agent workers exhibit stronger alignment with program-using human steps than those using non-programmatic ones. This highlights a defining characteristic of agent workers — a programmatic mode of operation. This is not only true for coding-oriented OpenHands agents, but also for general-purpose computer-use agents such as ChatGPT and Manus. While this programmatic tendency is understandable in tasks like data analysis — domains well represented in LMs trained on code data — this programmatic behavior notably persists even in visually intensive tasks such as slide creation and graphical design. Not only is such ‘blind’ visual creation possible, but LM-supported agents may in fact be more proficient at editing in the symbolic space (i.e., write programs) than in the visual space (i.e., adjust pixels) (Qiu et al., 2024).
@@ D.1 Which Domains Do Agents Have Advantages?
We compare completion time against work quality in Figure 19. While agents generally deliver work faster than human workers across most domains, we find that writing and data analysis are the areas where agent-produced work mostly approaches human-level quality. Unfortunately, for domains requiring less specialized expertise — such as administrative and repetitive computational tasks, where AI automation could potentially free up low-wage human labor — agents only do a mediocre job and are probably not ready for reliable deployment yet (Dell’Acqua et al., 2023). These findings suggest that further training is needed to strengthen fundamental capabilities such as visual understanding and UI interaction.
@@
$\bullet$ Computation  These tasks, typically considered basic and repetitive for human workers, primarily require basic visual skills (e.g., extracting numbers from bill images) and performing repetitive processes over a long horizon (e.g., thousands of steps). Unfortunately, agents currently lack robust visual understanding and long-horizon planning abilities, making them ill-suited to automate such work. This finding is somewhat disappointing, as these lower-entry-bar roles are among those most desirable for early automation. Instead, agents appear more capable in expert-oriented domains such as engineers (Cui et al., 2025), suggesting that these routine computational jobs may remain human-dominated for some time.
$\bullet$ Writing  While Figure 19 shows that agents are the closest at automating writing jobs currently done by humans, this may be partly due to the structured nature of the writing tasks in our study — HR drafting job descriptions or analysts composing financial reports with predefined modules. In contrast, many other forms of writing (e.g., books, articles) are considerably more open-ended and creative. Nevertheless, when viewed through the lens of writing as a skill rather than writing as a profession, many occupational writing activities (e.g., by engineers or analysts) tend to follow structured patterns. As prior studies in domain-specific writing, such as legal (Choi and Schwarcz, 2024) and academic (Lee et al., 2022; Zhao et al., 2025) contexts, report varying findings, our findings should be interpreted with caution and contextual awareness.
@@ D.2 Task Delegation by Programmability
$\bullet$ Less Programmable: Some tasks rely heavily on visual perception and lack a deterministic programmatic solution. For instance, in the task of viewing bill images and extracting data from them (Figure 21, right), a programmatic solution may require non-deterministic modules, such as invoking a neural OCR model to parse out the text, which may not guarantee task success (Wang et al., 2024a). Even human engineers, who can easily script some tasks such as data analysis, encounter challenges in this case. As in Figure 21 (right), the human worker’s programmatic attempt fails after searching for relevant OCR packages, observing undesired parsing performance, and ultimately abandoning the task. For such cases, it is often far easier for humans to simply inspect the images and manually enter the numbers in an Excel sheet via the UI. Although these tasks may be long-tail, particularly after more tasks become programmatically solvable, they remain inevitable components of computer-use activities. Many of our analysis provides substantial evidence of agents’ limitations in these less-programmable scenarios, underscoring the need for continued improvement through foundation model training.

### v2

@@ 7 Related Work
Agent Performance at Work  With the rise of computer-use agents, most efforts have focused on non-work settings such as online shopping (yao2022webshop; zhou2024webarena) and travel planning (deng2023mindweb). Beyond acting as personal assistants, agents hold growing potential to automate professional work and are projected to exert substantial labor market impacts (grace2018will; eloundou2023gpts). Several studies have examined agents’ autonomous performance in specific domains, including engineering (jimenez2024swebench), software services (drouin2024workarena), design (si2025design2code; ge2025autopresent), research (wijk2024re; lu2024ai; si2025can), and general business activities (wornow2024wonderbread; shen2024towards; huang2025crmarena); as well as their collaboration with humans in legal (choi2024ai; das2025lawflow) and writing (clark2018creative) contexts. TheAgentCompany benchmark (xu2024theagentcompany) advances this line of inquiry by evaluating multiple occupations, though it remains confined to the scope of a software engineering company. In this work, we seek to achieve a holistic coverage of all computer-related occupations, estimated to affect 71.9% of daily job functions, providing a more representative view of agent performance at work.
@@
Understanding Human Workflows  Prior research has examined human workflows in certain professional activities, such as customer service (brynjolfsson2025generative), design (son2024demystifying), and writing (dang2025corpusstudio). More recently, with the rise of AI agents, many studies have explored how human workflows change when AI tools are introduced into work. These studies observe a shift in human activities from ‘doing’ to ‘managing’, which tends to reduce cognitive load (shukla2025skilling), yet may also decrease overall productivity (shukla2025skilling; simkute2025ironies) and promote dishonest behaviors (kobis2025delegation).
@@
While these studies offer insights on how humans use and adapt to AI agents, most overlook the scenario in which AI agents automate human work entirely — arguably the most concerning case for the workforce. In such automation settings, human workflows differ substantially from those in augmentation scenarios (chen2025code). As employment for entry-level positions already faces threats (brynjolfsson2025canaries), a head-to-head comparative analysis between agents and humans — covering both behavior patterns and outcome quality — is crucial for identifying human advantages and informing policy and design decisions. Existing comparative efforts have limited contexts, such as ambiguity resolution in web tasks (son2024unveiling), a single web design task (loop112025ai), or secondary analysis on research articles (vaccaro2024combinations). To truly understand how agents as workers may reshape today’s human-centered workspace, we argue for a fine-grained, workflow-level investigation spanning multiple occupations and skill domains, to enable findings that shed light on diverse areas of work.
@@
Inducing Computer Workflows  Several tools have been developed for recording human trajectories (wang2025opencua; shaikh2025creating). However, such raw activity data are often hard to interpret, as low-level actions are task-agnostic and lack clear semantic grounding. Even for humans themselves, articulating explicit goals for certain actions, e.g., random clicks on blank areas driven by hesitation or anxiety, can be challenging. This poses challenges in conducting the head-to-head comparisons between human and agent computer-use activities. Inspired by how humans naturally communicate and coordinate around higher-level goals, we conceptualize and model activities at a coarser granularity through workflows, defined as sequences of actions frequently undertaken to achieve a specific goal (wang2025agent). Prior efforts toward this goal have largely depended on manual annotation (sodhi2023heap; grunde2025designing). In contrast, we introduce the first automated tool capable of inducing interpretable workflows from raw mouse and keyboard traces, generating structured goals at multiple levels of granularity to support diverse analytical needs. While we primarily employ this tool to study agent and human worker behaviors in this work, it also holds promise for broader applications, such as agent skill learning and human-agent collaboration.
@@ 8 Conclusion
The O*NET Database  The O*NET database provides a holistic view of occupational tasks across the U.S. labor market. However, the inherent difficulty in constructing and maintaining such a large-scale database may introduce inaccuracies. Although we utilized data from the most recent 2024 release, it may not fully reflect the current workforce distribution. Task requirements for certain occupations may have evolved, while others may have become obsolete or newly emerged in the current job market (acemoglu2022artificial).
Understanding AI’s Impact on Work  It is natural for many to wonder “will AI take my job?” especially amid the rapid advances in AI over the past few years. Beyond offering insights to researchers of AI agents and human-AI interactions, we hope this work also speaks to the broader community of human workers whose occupations may be affected by AI adoption. Because many in the general public may have little exposure to how AI actually functions, we aim for this study to serve as an accessible and detailed account that helps demystify their capabilities and limitations — encouraging informed awareness rather than ungrounded fear or hype. While parts of our study may reveal the uncomfortable reality that agents may gradually take over fixed-procedure tasks as they advance, many essential aspects of human work, such as communication and decision-making, still demand substantial technological progress and regulatory considerations before comparable automation becomes feasible (comunale2024Economic). Overall, our goal is to offer a clearer, evidence-based understanding of how these seemingly powerful AI agents operate and to inform a grounded, proactive perspective on how humans and AI can move forward together.
@@ Acknowledgments
We would like to thank Sherry Tongshuang Wu, Yuxuan Li, Christina Qianou Ma, Zhoujun Cheng, Xuhui Zhou, Yuxiao Qu, Lawrence Jang, Xinran Zhao, Saujas Vaduguru, Vijay Viswanathan, Chenyang Yang, Mingqian Zheng, Youssouf Kebe, John Yang, Vishakh Padmakumar, Jiaxin Pei, and Alexander Spangher for their helpful discussions and insightful feedback on the project. Zora Zhiruo Wang is supported by Google PhD Fellowship. This research is supported in part by grants from Sloan Foundation, ONR grant N000142412532 and Open Philanthropy.
@@ A.3 Human Activity Collection
Processing Raw Computer-Use Activities  Computer-use activities collected from human real-world usage can be redundant, due to the limitations in the recording tool implementation. For example, each mouse scrolling (scroll()) and keyboard press (keypress(‘h’)) event is recorded as an independent action. These human actions execute with arguments that are more fine-grained (e.g., keypress(‘h’), keypress(‘o’), $\cdots$, keypress(‘?’)) than arguments in the agent actions (e.g., keypress(‘how ai agents do human work?’)), thus becoming costly to process and hard to interpret. We hence perform post-processing to merge consecutive keypress and scroll actions into single actions with concatenated arguments. We also identify possible double clicks if two click() actions happen within 0.1 seconds; and merge them into a single double_click() action. This effectively simplifies the trajectory data, reducing the action count by 83.2% (from 5831.1 to 981.1 actions) in an average human trajectory, meanwhile aligning better with the granularity of agent computer activities (wang2025opencua).
@@ A.4 Agent Activity Collection
While we can directly access and store all actions and states for OpenHands agents, we can only access the information presented at the ChatGPT and Manus agent UI interfaces. For each step in the ChatGPT and Manus trajectories, the agent will elaborate its thought process in text, and present a screenshot of the virtual computer screen. We infer the actions taken from the agent’s thoughts by mapping back to the common computer-use agent action space such as zhou2024webarena; xie2024osworld. We present the actions employed by all agent frameworks in Table 4. Nonetheless, action names for ChatGPT and Manus may not be precisely what was executed by the agent, but are our best guesses and should serve a similar functionality to the actual action, which we unfortunately do not have access to. Meanwhile, the listed actions are only actions that we observe agents have taken while completing all the task instances in our study. These actions may not constitute the full action space available to the agents.
@@ C.1 Do Agents Follow Human Task Workflows?
As shown in Table 7, agent workers exhibit stronger alignment with program-using human steps than those using non-programmatic ones. This highlights a defining characteristic of agent workers — a programmatic mode of operation. This is not only true for coding-oriented OpenHands agents, but also for general-purpose computer-use agents such as ChatGPT and Manus. While this programmatic tendency is understandable in tasks like data analysis — domains well represented in LMs trained on code data — this programmatic behavior notably persists even in visually intensive tasks such as slide creation and graphical design. Not only is such ‘blind’ visual creation possible, but LM-supported agents may in fact be more proficient at editing in the symbolic space (i.e., write programs) than in the visual space (i.e., adjust pixels) (qiu2024can).
@@ D.1 Which Domains Do Agents Have Advantages?
We compare completion time against work quality in Figure 19. While agents generally deliver work faster than human workers across most domains, we find that writing and data analysis are the areas where agent-produced work mostly approaches human-level quality. Unfortunately, for domains requiring less specialized expertise — such as administrative and repetitive computational tasks, where AI automation could potentially free up low-wage human labor — agents only do a mediocre job and are probably not ready for reliable deployment yet (dell2023navigating). These findings suggest that further training is needed to strengthen fundamental capabilities such as visual understanding and UI interaction.
@@
$\bullet$ Computation  These tasks, typically considered basic and repetitive for human workers, primarily require basic visual skills (e.g., extracting numbers from bill images) and performing repetitive processes over a long horizon (e.g., thousands of steps). Unfortunately, agents currently lack robust visual understanding and long-horizon planning abilities, making them ill-suited to automate such work. This finding is somewhat disappointing, as these lower-entry-bar roles are among those most desirable for early automation. Instead, agents appear more capable in expert-oriented domains such as engineers (cui2025effects), suggesting that these routine computational jobs may remain human-dominated for some time.
$\bullet$ Writing  While Figure 19 shows that agents are the closest at automating writing jobs currently done by humans, this may be partly due to the structured nature of the writing tasks in our study — HR drafting job descriptions or analysts composing financial reports with predefined modules. In contrast, many other forms of writing (e.g., books, articles) are considerably more open-ended and creative. Nevertheless, when viewed through the lens of writing as a skill rather than writing as a profession, many occupational writing activities (e.g., by engineers or analysts) tend to follow structured patterns. As prior studies in domain-specific writing, such as legal (choi2024ai) and academic (lee2022coauthor; zhao2025knoll) contexts, report varying findings, our findings should be interpreted with caution and contextual awareness.
@@ D.2 Task Delegation by Programmability
$\bullet$ Less Programmable: Some tasks rely heavily on visual perception and lack a deterministic programmatic solution. For instance, in the task of viewing bill images and extracting data from them (Figure 21, right), a programmatic solution may require non-deterministic modules, such as invoking a neural OCR model to parse out the text, which may not guarantee task success (wang2024what). Even human engineers, who can easily script some tasks such as data analysis, encounter challenges in this case. As in Figure 21 (right), the human worker’s programmatic attempt fails after searching for relevant OCR packages, observing undesired parsing performance, and ultimately abandoning the task. For such cases, it is often far easier for humans to simply inspect the images and manually enter the numbers in an Excel sheet via the UI. Although these tasks may be long-tail, particularly after more tasks become programmatically solvable, they remain inevitable components of computer-use activities. Many of our analysis provides substantial evidence of agents’ limitations in these less-programmable scenarios, underscoring the need for continued improvement through foundation model training.